OpenStudy (anonymous):

Solve the system of equations. { 3x-3y= 5 {-12x+12y=-15

hero (hero):

3x - 3y = 5 -12x + 12y = -15 ---> -3(4x - 4y = -3(5) ---> 4x - 4y = 5 3x - 3y = 5 4x - 4y = 5 y = (5 - 3x)/-3 y = (5 - 4x)/-4 y = y -4(5 - 3x) = -3(5 - 4x) -20 + 12x = -15 + 12x 0 = 5 false

hero (hero):

No solution